OpenStudy (anonymous):

25+144 squared = 13 correct?

OpenStudy (anonymous):

If you meant 13^2, yes, that's a Pythagorean triple in which 5^2 + 12^2 = 13^2 You have your numbers a little bit mixed up.

OpenStudy (anonymous):

no i dont mean that

OpenStudy (anonymous):

Then no, of course not. 25 + 144^2 = 13? 25 by itself does not equal 13, much less if you add 20,736, which is 144^2

OpenStudy (anonymous):

Wait, do you mean this: \[\sqrt(25+144)\]

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

Ahh ok, that's not called "squared" that's called square ROOT haha if you say square you multipy a number by itself lol. Ok, then yes, your first statement was right. When you add 25 to 144 you get 169. The Square Root of 169 is 13, haha lol

OpenStudy (anonymous):

oh thanks man